1.The two fair dice: black and red, with 6 sideS are rolled. WRITE A SAMPLE SPACE FOR AN EVENT DESCRIBED BELOW:“The sum of face up numbers 5”.

2. What is the probability that the sum of dace up numbers 5?
Answer #1

here let (a,b) is the outcome on a single roll where a is outcome on black dice and b is outcome on red dice,

hence sample space for event " The sum of face up numbers 5 " ={(1,4),(2,3),(3,2),(4,1)}

2)

total number of events on rolls of two dice =6*6 =36

number of events so that sum of face up numbers 5 =4

hence probability that the sum of dace up numbers 5 =4/36 =1/9